1. The Bok Chicken Factory is trying to figure out how to minimize the cost of producing 1200 units of chicken parts. The production function is q = 100L0.5 K0.5. The wage rate is birr 9 per hour and the rental rate on capital is birr 4 per machine hour.




A. Find the minimum cost of producing 1200 units.




B. Find the maximum output that can be produced for a total cost of birr 720.





1
Expert's answer
2022-01-23T15:39:48-0500

Solution:

A.). Q = 100L0.5K0.5

Derive MPL and MPK

MPL = "\frac{\partial Q} {\partial L}" = 50L-0.5K0.5


MPK = "\frac{\partial Q} {\partial K}" = 50L0.5K-0.5


"\frac{MPL}{MPK}" = "\frac{w}{r}"

w = 9

r = 4

50L-0.5K0.5 "\div" 50L0.5K-0.5 = "\frac{9}{4}"

K = "\frac{9L}{4}" = 2.25L


1200 = 100L0.5K0.5

1200 = 100L0.5(2.25L0.5)

L = 5.3

K = 2.25L = 2.25(5.3) = 11.925

C = wL + rK

C = (9 "\times" 5.3) + (4 "\times" 11.925) = 47.7 + 47.7 = 95.4

The minimum cost to produce 1200 units = 95.4